OpenStudy (anonymous):

Solve each sysytem. In your work, identify the shape of the graph of each equation. 1.) {7y^2+x^2=64 x+y+4 2.) {x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

OpenStudy (ryamorgan284):

Use this site https://mathway.com/

OpenStudy (anonymous):

I don't think it does this kinda question but thanks

OpenStudy (ryamorgan284):

it will trust me...just click the link and type in tour problem..

OpenStudy (mathmale):

You'll need to learn the general equations for geometric shapes such as circle, parabola, ellipse, hyperbola, and so on. Look up "equations of conic sections." Try this, and then post some questions of your own, so that someone can help you get on track. Note that x+y=4 is a linear function; its graph is a straight line. The other functions, the ones with x^2 and y^2 in them, are curves. It's your job to determine the names and shapes of these curves.

OpenStudy (anonymous):

Alright thanks!

OpenStudy (alivejeremy):

need help

OpenStudy (anonymous):

Yes

OpenStudy (alivejeremy):

five years later but ok

OpenStudy (alivejeremy):

so do u post to work it out or something

OpenStudy (alivejeremy):

im not sure :)

OpenStudy (anonymous):

Yea

OpenStudy (alivejeremy):

is this like terms

OpenStudy (alivejeremy):

TAG ME WHEN U GET BACK

OpenStudy (alivejeremy):

BYE

OpenStudy (anonymous):

I'm here... @alivejeremy

OpenStudy (anonymous):

All it says is to solve the system

OpenStudy (alivejeremy):

{7y^2+x^2=64x+y+4

OpenStudy (alivejeremy):

it post to look like that

OpenStudy (alivejeremy):

or nah

OpenStudy (anonymous):



OpenStudy (anonymous):

Like that, just 1 big bracket

OpenStudy (alivejeremy):

oh kool

OpenStudy (anonymous):

It's ok if you can't help

OpenStudy (alivejeremy):

i can im working it out

OpenStudy (alivejeremy):

whats the name of this math tho

OpenStudy (anonymous):

pre-cal

OpenStudy (alivejeremy):

oh kool

OpenStudy (alivejeremy):

look at that

OpenStudy (anonymous):

Ok?

OpenStudy (alivejeremy):

u get it

OpenStudy (alivejeremy):

thats number 1

OpenStudy (alivejeremy):

number 1. 16 + y^2 - 8y

OpenStudy (alivejeremy):

16 + y^2 - 8y

OpenStudy (triciaal):

do you still need help?

OpenStudy (anonymous):

Yes, with the second one

OpenStudy (triciaal):

{x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

OpenStudy (triciaal):

suggestion complete the square for each equation then set them equal

OpenStudy (alivejeremy):

x^2+y^2+2x+2y=0 |completing squares x^2+2x+ y^2+2y =0 (x+1)^2 -1 + (y+1)^2 -1 = 0 (x+1)^2 + (y+1)^2 = 2 x^2+y^2+4x+6y+12=0 |completing squares x^2+4x + y^2+6y +12=0 (x+2)^2 -4 + (y+3)^2 -9 + 12 = 0 (x+2)^2 + (y+3)^2 -9 = 1 Algebraically: x^2+y^2+2x+2y = x^2+y^2+4x+6y+12

OpenStudy (triciaal):